OpenStudy (anonymous):

help me:(((( i will give medalssss Daniel is looking over some data regarding the temperatures of countries and how that relates to their number of mosquitos. The data seem to be exponential. How can Daniel determine if the data are exponential, and how can he convert that to a logarithmic function?

OpenStudy (anonymous):

@amistre64 @phi

OpenStudy (phi):

make a table of temperature versus log(# of skeeters) see if you can draw a straight line through this new data.

OpenStudy (anonymous):

Would I just make up numbers? I wasn't given any… I think its just asking how to know if data is exponential and how to convert it into log.

OpenStudy (phi):

exponential means that to get to the next number, you multiply by some factor "r" for example, if r=2, then the numbers 1,2,4,8,16 is exponential if you take the ratio of any pair, you will get r. example: 2/1 = 2 or 16/8 = 2 that is one way to test if the numbers are exponential

OpenStudy (phi):

if you are able to find "r" that works, then you can write the equation as \[ y = A r^x \] to find A, find the value of your numbers for x=0 for example, if we have the points (0,1), (1,2), (2,4), (3,8), (4,8) r is 2 (ratio of two adjacent "y" values) using the point (0,1) (where x=0 and y=1) in the equation y = A r^x 1 = A 2^0 2^0 is 1, so 1 = A*1 or 1=A A is 1, so the equation is \[ y = 2^x \]

OpenStudy (anonymous):

y=2